$SPAD/src/input macros.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool macros.output
)set message test on
)set message auto off
)clear all
--S 1 of 4
macro I == Integer
--R
--R                                                                   Type: Void
--E 1
--S 2 of 4
macro M(R) == Matrix(R)
--R
--R                                                                   Type: Void
--E 2
--S 3 of 4
macro p(n) == x < n
--R
--R                                                                   Type: Void
--E 3
--S 4 of 4
macro q(i,j) == if x < i then i else j
--R
--R                                                                   Type: Void
--E 4
)spool
)lisp (bye)

———-

References^{[1] nothing}

4